# Attention Mechanisms Don’t Learn Additive Models: Rethinking Feature Importance for Transformers

### v1

@@ F.7 Runtime analysis
(a) BERT
Runtime (s) SHAP LIME IG Grad SLALOM
1000 Samples 3.89 $\pm$ 0.21 1.55 $\pm$ 0.22 0.05 $\pm$ 0.01 0.01 $\pm$ 0.01 0.24 $\pm$ 0.02
3000 Samples 13.70 $\pm$ 0.29 4.69 $\pm$ 0.66 0.05 $\pm$ 0.01 0.01 $\pm$ 0.01 0.71 $\pm$ 0.01
10000 Samples 45.54 $\pm$ 0.80 15.57 $\pm$ 2.20 0.05 $\pm$ 0.01 0.01 $\pm$ 0.01 2.40 $\pm$ 0.02
@@
Runtime (s) SHAP LIME IG Grad SLALOM
@@
While IG and Gradient explanations are the quickest, they also require backward passes which have large memory requirements. As expected, the computational complexity for surrogate model explanation (LIME, SHAP, SLALOM) is dominated by the number of samples and forward passes done. Our implementation of SLALOM is around 8-10x faster than LIME and almost 20x faster than SHAP (all approaches used a GPU-based, batching-enabled implementation), which we attribute to the fact that SLALOM can be fitted using substantially shorter sequences than are used by LIME and SHAP.
We are interested to find out how many samples are required to obtain an explanation of comparable quality to SHAP. We successively increase the number of samples used to fit our surrogates and report the performance in the deletion benchmark (where the prediction should drop quickly when removing the most important tokens). We report the Area over the Perturbation Curve (AOPC) as before (this corresponds to their Comprehensiveness metric of ERASER[16], higher scores are better). We compare the performance to shap.KernelExplainer.shap_values(nsamples=auto) method of the shap package in Table 12.
@@
F.8 Applying SLALOM to Large Language Models
@@
Our work is mainly concerned with sequence classification. In this application, we observe mid-sized models like BERT to be prevalent. On the huggingface hub, among the 10 most downloaded models on huggingface, 9 are BERT-based and the remaining one is another transformer with around 33M parameters [^2] https://huggingface.co/models?pipeline_tag=text-classification&sort=downloads In common benchmarks like DBPedia classification [^3] https://paperswithcode.com/sota/text-classification-on-dbpedia, the top-three models are transformers with two of them also being variants of BERT. We chose our experimental setup to reflect this. Nevertheless, we are interested to see if SLALOM can provide useful insights for larger models as well and therefore experiment with larger models. To this end, we use a model from the BLOOM family [30] with 7.1B parameters as well as the recent Mamba model (2.8B) [21] on the Yelp-HAT dataset and compute SLALOM explanations. Note that the Mamba model does not even follow the transformer framework considered in this work. We otherwise follow the setup described in Figure 3 and assess whether our explanations can predict human attention. The results in Table 13 highlight that this is indeed the case, even for larger models. The ROC scores are in a range comparable to the ones obtained for the smaller models. For the non-transformer Mamba model we observe a drop in the value of the importance scores. This may suggest that value scores and linearized SLALOM scores are more reliable for large, non-transformer models.
LM values $v$ importances $s$ lin. $(\exp(s)\cdot v)$
BLOOM-7B 0.69 $\pm$ 0.01 0.71 $\pm$ 0.02 0.70 $\pm$ 0.02
Mamba-2.8B 0.69 $\pm$ 0.02 0.32 $\pm$ 0.02 0.70 $\pm$ 0.01
Table 13: ROC-Scores for predicting Human attention with SLALOM using LLMs

### v2

@@ F.6 Runtime analysis
(a) DistilBERT
Approach / # samples 1000 2000 5000 10000
SHAP 2.35 $\pm$ 0.01 4.62 $\pm$ 0.02 11.56 $\pm$ 0.03 23.08 $\pm$ 0.08
LIME 0.80 $\pm$ 0.04 1.58 $\pm$ 0.07 3.93 $\pm$ 0.19 8.04 $\pm$ 0.39
SLALOM-fidel 0.74 $\pm$ 0.03 1.42 $\pm$ 0.06 3.77 $\pm$ 0.24 7.95 $\pm$ 0.41
SLALOM-eff 0.42 $\pm$ 0.01 0.80 $\pm$ 0.01 2.03 $\pm$ 0.01 4.13 $\pm$ 0.02
LRP 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00
IG 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00
Grad 0.01 $\pm$ 0.00 0.01 $\pm$ 0.00 0.01 $\pm$ 0.00 0.01 $\pm$ 0.00
@@
Runtime (s) SHAP LIME IG Grad SLALOM-eff
@@
While IG and Gradient explanations are the quickest, they also require backward passes which have large memory requirements. As expected, the computational complexity for surrogate model explanation (LIME, SHAP, SLALOM) is dominated by the number of samples and forward passes done. Our implementation of SLALOM is around 2x faster than LIME and almost 5x faster than SHAP (all approaches used a GPU-based, batching-enabled implementation), which we attribute to the fact that SLALOM can be fitted using substantially shorter sequences than are used by LIME and SHAP.
We are interested to find out how many samples are required to obtain an explanation of comparable quality to SHAP. We successively increase the number of samples used to fit our surrogates and report the performance in the deletion benchmark (where the prediction should drop quickly when removing the most important tokens). We report the Area over the Perturbation Curve (AOPC) as before (this corresponds to their Comprehensiveness metric of ERASER (DeYoung et al., 2020), higher scores are better). We compare the performance to shap.KernelExplainer.shap_values(nsamples=auto) method of the shap package in Table 15.
@@
F.7 Applying SLALOM to Large Language Models
@@
Our work is mainly concerned with sequence classification. In this application, we observe mid-sized models like BERT to be prevalent. On the huggingface hub, among the 10 most downloaded models on huggingface, 9 are BERT-based and the remaining one is another transformer with around 33M parameters [^3] https://huggingface.co/models?pipeline_tag=text-classification&sort=downloads (as of September 2023). In common benchmarks like DBPedia classification [^4] https://paperswithcode.com/sota/text-classification-on-dbpedia, the top-three models are transformers with two of them also being variants of BERT. We chose our experimental setup to reflect this. Nevertheless, we are interested to see if SLALOM can provide useful insights for larger models as well and therefore experiment with larger models. To this end, we use a model from the BLOOM family (Le Scao et al., 2023) with 7.1B parameters as well as the recent Mamba model (2.8B) (Gu & Dao, 2023) on the Yelp-HAT dataset and compute SLALOM explanations. Note that the Mamba model does not follow the transformer framework considered in this work. We otherwise follow the setup described in Figure 3 and assess whether our explanations can predict human attention. The results on the bottom of Table 2(b) highlight that this is indeed the case, even for larger models. The ROC scores are in a range comparable to the ones obtained for the smaller models. For the non-transformer Mamba model we observe a drop in the value of the importance scores. This may suggest that value scores and linearized SLALOM scores are more reliable for large, non-transformer models.
Applying SLALOM to blackbox models. Finally, we would like to emphasize that SLALOM, as a surrogate model can be applied to black-box models as well. To impressively showcase this, we apply SLALOM to OpenAI’s GPT-4 models via the API (Appendix F.7). We use the larger GPT-4-turbo and smaller GPT-4o-mini for comparison. We prompt the model with the following template to classify the review and only output either 0 or 1 as response.
SYSTEM: You are assessing movie reviews in an online forum. Your goal is to assess the reviews’ overall sentiment as ’overall negative’ (label ’0’) or ’overall positive’ (label ’1’). Your will see a review now and you will output a label. Make sure to only answer with either ’0’ or ’1’. USER: $<$the review$>$
We then use the token probabilities returned by the API to compute the score. We create 500 samples as a training dataset for the surrogate model and fit SLALOM the model using SLALOM-eff. We obtain the importance plots shown in Figure 13. This highlight that SLALOM scales up to large models. However, we would like to stress that there can be no guarantees as we have no knowledge about the specific structure of the model.
(a) GPT-4-turbo
(b) GPT-4o-mini
Figure 13: We apply SLALOM to OpenAI GPT-4-turbo and GPT-4o to showcase its scalability. We use SLALOM-eff scores and linearize them to obtain the highlighed attributions.
F.8 Case study: Identifying Vulnerabilites and Weaknesses in classification models with SLALOM
In our public code repository, we provide a case study to highlight how SLALOM can be used to uncover practical flaws in transformer classification models. We use a trained BERT model on IMDB sentiment classification as an example and compute SLALOM scores for the tokens in the first 15 samples of the test set ($\approx$ 1200 tokens). SLALOM’s scores make it easy to see which tokens have a potentially severe impact on the movie review outcome. We first select all tokens with an importance score > 2, all other words have no substantial impact, in particular when added to a larger sequence. We provide a visualization of these tokens in Figure 14. We then make some interesting discoveries mainly based on value scores:
  1. 1.
    There are more words that are negatively interpreted by the model than positive words
  2. 2.
    Out of the words that have highly negative value scores (+importance >2), we identify several words that are some that are not directly negatively connotated, e.g., “anyway”, “somehow”, “never”, “anymore”, “probably”, “doesn”, “maybe”, “without”, “however”, “surprised”
We then show that by a few (4) minor modification steps, e.g., adding some of these words to a review, we can change the classification decision for a review from positive to negative, without essentially altering its content (i.e., we manually construct an adversarial example.) This highlights how SLALOM can intuitively help to uncover 1) the influential tokens that contribute most to the decision and 2) allow for a fine-grained analysis that can help uncover spurious concepts and give practitioners an intuitive understanding of a model’s weaknesses.
Figure 14: High-impact tokens for BERT identified in our case study. Use zoom for best view.